How much heat is required to change 200ml of ice at - 22°C into steam 100 °C


1
Expert's answer
2021-04-26T17:09:17-0400

"\lambda=3.4\cdot 10^5~\frac{J}{kg},\\
c_i=2.1\cdot 10^3~\frac{J}{kg\cdot °C},\\
c_w=4.2\cdot 10^3~\frac{J}{kg\cdot °C},\\
L=2.26\cdot 10^6~\frac{J}{kg},"

"\Delta t_1=0-(-22)=22°C,\\
\Delta t_2=100-0=100°C,"

"m=\rho V."

"Q=Q_1+Q_2+Q_3+Q_4=\rho V(\lambda+c_i \Delta t_1+c_w \Delta t_2+L),"

"Q=613~kJ."